# Probabilistic Dynamic Line Rating Forecasting with Line Graph Convolutional LSTM

### v1

@@ I Introduction
    We propose a novel network-wide probabilistic DLR forecasting framework called double-hop line graph convolutional network (D-LGCLSTM) that combines a double-hop line graph convolutional network with LSTM to effectively capture complex spatio-temporal patterns in transmission networks. By utilizing double-hop message passing, D-LGCLSTM captures extended spatial correlations and reduces feature duplication within a single layer. To the best of our knowledge, this is the first work that provides probabilistic DLR forecasting by incorporating both spatial and temporal information across entire transmission networks.
@@
    We find that the forecasting performance of the single-hop line graph convolutional network (hereafter referred to as LGCLSTM) is degraded due to feature duplication, where similar inputs are repeatedly aggregated during the message-passing process. We show that the proposed D-LGCLSTM can effectively mitigate the adversarial effect of feature duplication and capture extended spatial patterns across the network while using 65% fewer parameters compared to LGCLSTM.
@@
    We rigorously evaluate D-LGCLSTM against three state-of-the-art algorithms and LGCLSTM in probabilistic DLR forecasting [3, 12, 13] on the Texas 123-bus backbone transmission system using five years of historical data. We extensively demonstrate that D-LGCLSTM outperforms all the baselines in terms of reliability, sharpness, and the number of learnable parameters.
@@ III-A1 Data Preparation
We use the Texas 123-bus backbone transmission (TX-123BT) system [16] to verify the performance of the proposed method in probabilistic DLR forecasting. This system contains 123 buses and 244 lines. For DLR forecasting, we reduce the number of lines to 173 by merging parallel lines. We utilize five years of historical weather data for each bus and DLR data for each line from January 1, 2017, to December 31, 2021, with a one-hour resolution. The weather data include measurements of temperature, wind speed, wind direction, and solar radiation. The DLR data consists of line ratings calculated based on heat balance equation [17]. We split the dataset into a training set and a testing set using a 4:1 ratio. Each bus includes the previous seven days of historical weather data and its geographical coordinates. Each line includes the previous seven days of historical DLR data, its length, and the current season (spring, summer, fall, or winter). The model uses these input data to forecast the next day’s DLR for each line.
@@
Figure 4: Heat maps of the average QS for each transmission line using test data across TX-123BT system. The Gray dotted arrow points the line 123 where the highest rate of change in QS from LSTM to D-LGCLSTM is observed.
@@ III-A3 Baseline Models
We compare the proposed D-LGCLSTM against four baselines in Table I. LSTM [3] captures only temporal patterns of a single line. T-GCN [12] combines GCN and LSTM sequentially but does not integrate the GCN into the LSTM cell. In contrast, GCLSTM [13] integrates the GCN directly into the LSTM cell. Both T-GCN and GCLSTM operate on the original graph without transforming it into a line graph. LGCLSTM applies GCLSTM after line graph conversion for consistent feature dimension. D-LGCLSTM advances further by aggregating the features over double-hop neighbors in the line graph.
@@ III-B1 Overall Performance Comparisons
Table II provides a comparison of DLR forecasting performance across the baselines. The proposed D-LGCLSTM consistently outperforms the baseline models across all evaluation metrics. Specifically, D-LGCLSTM achieves nearly half the ACE of LSTM. In addition, LSTM shows higher ACE compared to all the models in Table II, which demonstrates the necessity of using graph-based model for reliable forecasting. Furthermore, D-LGCLSTM attains a significantly lower PINAW and thus obtains sharper prediction intervals compared to T-GCN, which does not integrate GCN or LGCN into LSTM.
Moreover, D-LGCLSTM outperforms all the baselines in IS and QS, which measure both reliability and sharpness. D-LGCLSTM reduces nearly 7% of IS and QS from GCLSTM by applying a double-hop message passing in a line graph. Thus, D-LGCLSTM successfully achieves high reliability while keeping the prediction intervals as sharp as possible. This is highly beneficial from a power system perspective, as it enables operators to make more informed and precise decisions by maximizing transmission line utilization without unnecessary conservatism.
@@ IV Conclusion
In this paper, we proposed a novel network-wide probabilistic dynamic line rating (DLR) forecasting model called double-hop line graph convolutional LSTM (D-LGCLSTM), which integrates line graph convolutional networks into LSTM to incorporate both spatial and temporal information. By employing a double-hop message passing on the line graph, D-LGCLSTM captures extended spatial correlations and mitigates feature duplication in single-hop models. The simulations on the Texas 123-bus backbone transmission system demonstrate that D-LGCLSTM outperforms all the baselines in terms of reliability and sharpness while using the least number of parameters. Specifically, D-LGCLSTM achieves up to a 7% improvement in IS and QS and reduces the number of model parameters by at most 99% compared to baselines. For future work, we plan to integrate D-LGCLSTM with grid operations, such as security-constrained unit commitment or market operations, to further analyze its impact on the power systems.

### v2

@@ I Introduction
    We propose a novel network-wide probabilistic DLR forecasting framework called double-hop line graph convolutional LSTM (D-LGCLSTM) that combines a double-hop line graph convolutional network with LSTM to effectively capture complex spatio-temporal patterns in transmission networks. By utilizing double-hop message passing, D-LGCLSTM captures extended spatial correlations and reduces feature duplication within a single layer. To the best of our knowledge, this is the first work that provides probabilistic DLR forecasting by incorporating both spatial and temporal information across entire transmission networks.
@@
    We find that the forecasting performance of the single-hop line graph convolutional LSTM (hereafter referred to as LGCLSTM) is degraded due to feature duplication, where similar inputs are repeatedly aggregated during the message-passing process. We show that the proposed D-LGCLSTM can effectively mitigate the adversarial effect of feature duplication and capture extended spatial patterns across the network while using 65% fewer parameters compared to LGCLSTM.
@@
    We rigorously evaluate D-LGCLSTM against four state-of-the-art algorithms and LGCLSTM in probabilistic DLR forecasting [12, 3, 13, 14] on the Texas 123-bus backbone transmission system using five years of historical data. We extensively demonstrate that D-LGCLSTM outperforms all the baselines in terms of reliability, sharpness, and the number of learnable parameters.
@@ III-A1 Data Preparation
We use the Texas 123-bus backbone transmission (TX-123BT) system [17] to verify the performance of the proposed method in probabilistic DLR forecasting. This system contains 123 buses and 244 lines. For our experiments, we reduce the lines to 173 by retaining only one representative among parallel lines that have identical parameters (e.g., endpoints, weather, and rating data) to avoid redundant inputs and reducing computational overhead. We utilize five years of historical weather data for each bus and DLR data for each line from January 1, 2017, to December 31, 2021, with a one-hour resolution. The weather data include measurements of temperature, wind speed, wind direction, and solar radiation. The DLR data consist of line ratings calculated based on the heat balance equation based on IEEE 738 standard [18] as done in [7]. We split the dataset into a training set and a testing set using a 4:1 ratio, where the 2021 test data notably encompasses the 2021 Texas power crisis. Each bus includes the previous seven days of historical weather data and its geographical coordinates. Each line includes the previous seven days of historical DLR data, its length, and the current season (spring, summer, fall, or winter). The model uses these input data to forecast the next day’s DLR for each line. Note that all the experiments are conducted on NVIDIA A100 SXM4 40 GB GPU and Intel Xeon 2.20 GHz CPU.
@@ III-A3 Baseline Models
We compare the proposed D-LGCLSTM against five baselines in Table I. LSTM [3] captures only temporal patterns of a single line. T-GCN [13] combines GCN and LSTM sequentially but does not integrate the GCN into the LSTM cell. In contrast, GCLSTM [14] integrates the GCN directly into the LSTM cell. Both T-GCN and GCLSTM operate on the original graph without transforming it into a line graph. LGCLSTM applies GCLSTM after line graph conversion for consistent feature dimension. D-LGCLSTM advances further by aggregating the features over double-hop neighbors in the line graph. Note that since D-LGCLSTM leverages a single layer of double-hop LGCN, it requires fewer parameters and offers faster inference compared to GCLSTM and LGCLSTM, which rely on multiple layers. We also include Quantile Regression Forests (QRF) [12] as a non-deep-learning reference method that forecasts single-line ratings without leveraging graph structure. We train LSTM, GCLSTM, LGCLSTM, and D-LGCLSTM for 50 epochs, and T-GCN for 20 epochs using AdamW (learning rate and weight decay of $10^{-3}$). All models use an LSTM hidden dimension of 64, with a mini-batch size of 64 for T-GCN, GCLSTM, and LGCLSTM, and 128 for D-LGCLSTM.
@@ III-B1 Overall Performance Comparisons
Table II provides a comparison of DLR forecasting performance across the baselines. The proposed D-LGCLSTM outperforms baseline models in all evaluation metrics. Specifically, D-LGCLSTM achieves nearly half the ACE of LSTM. In addition, LSTM shows a higher ACE compared to all the models in Table II, demonstrating the necessity to use a graph-based model for reliable forecasting. Although QRF achieves the lowest ACE among all methods, it exhibits the worst PINAW. By contrast, D-LGCLSTM achieves a significantly lower PINAW and thus obtains sharper prediction intervals compared to T-GCN, which does not integrate GCN or LGCN into LSTM.
Moreover, D-LGCLSTM outperforms all the baselines in IS and QS, which measure both reliability and sharpness. By contrast, QRF achieves the worst IS and QS scores due to its unbalanced performance, which implies that simply minimizing ACE does not guarantee practical intervals. D-LGCLSTM reduces nearly 7% of IS and QS from GCLSTM by applying a double-hop message passing in a line graph. Thus, D-LGCLSTM successfully achieves high reliability while keeping the prediction intervals as sharp as possible. This is highly beneficial from a power system perspective, as it enables operators to make more informed and precise decisions by maximizing transmission line utilization without unnecessary conservatism.
Figure 4: Heat maps of the average QS for the lines in TX-123BT.
@@ IV Conclusion
In this paper, we proposed a novel network-wide probabilistic dynamic line rating (DLR) forecasting model called double-hop line graph convolutional LSTM (D-LGCLSTM), which integrates line graph convolutional networks into LSTM to incorporate both spatial and temporal information. By employing a double-hop message passing on the line graph, D-LGCLSTM captures extended spatial correlations and mitigates feature duplication in single-hop models. The simulations on the Texas 123-bus backbone transmission system demonstrate that D-LGCLSTM outperforms all the baselines in terms of reliability and sharpness while using the least number of parameters. Specifically, D-LGCLSTM achieves up to a 7% improvement in IS and QS and reduces the number of model parameters by at most 99% compared to baselines.
Although D-LGCLSTM effectively captures true DLR across the network better than the baselines, abrupt or extreme changes in DLR remain challenging. For future work, we plan to integrate D-LGCLSTM with grid operations, such as security-constrained unit commitment or market operations, including applications to larger or more diverse power networks, and investigate how operators can select suitable quantiles for real-world operational needs. We also plan to conduct a more detailed investigation of the model’s performance under extreme weather events to ensure reliable DLR forecasting.